OpenStudy (anonymous):

in need of some quick help! Assume that the variable x represents a positive real number Radical 18x^10 simplify

OpenStudy (anonymous):

\[\sqrt{18x^{10}}=\sqrt{3\times 3\times 2\times x^5\times x^5}\\= 3x^5\sqrt{2}\\=3\sqrt{2}x^5 \ [simplified \ form,\ can't\ further\ simplified]\]

OpenStudy (anonymous):

3 radical 2 x^5

OpenStudy (anonymous):

okay thank you! thats what i got too!

OpenStudy (anonymous):

cheers! well done!